2. Find the most general antiderivative.
a) ?(?) = 2?^2 -?+3 /?
b) ?(?) = √? − ???(?)???(?)

2. Find the most general antiderivative. b) g(t) = Vt - csc(t)cot(t) a) f(x) = 2x²-x+3
Answer #1

x @ fcal = 2x² - x + 3 = 2x-1 to fw = (x2-x +3868 +C) By fw) (F (60) TF (a)=x²-x+ 3 lux + c/ This is the required Anti deriva